OpenStudy (anonymous):

Write the equation in standard form (5/6, -4) and (3/4, -4)

OpenStudy (anonymous):

In the straight line equation: y = mx + c: 'x' and 'y' are the coordinates of the points that satisfy the function and so lie on the straight line graph. 'm' is the gradient of the straight line graph, and 'c' is the 'y intercept' of the straight line graph.

OpenStudy (anonymous):

Probably y=-4. Note that there is no change on the y value, so there is a zero slope.

OpenStudy (anonymous):

I nneed the equation in standard form, which is y=mx+b? Correct?

Directrix (directrix):

No. Standard Form of a Linear Equation: The standard form of a linear equation is Ax + By = C, where A, B, and C are integers, A0, and A and B are not both zero. ----- If y = -4 is the correct equation, then in standard form, 0x + (1) y = -4.

OpenStudy (anonymous):

I don't think a zero coefficient is usually included just to complete the standard form.

Directrix (directrix):

If the zero were not included, there would be no x-term in the standard form equation of the line. Therefore, the equation would not be in standard form. View more info at the link - scroll about halfway down the page. http://jwilson.coe.uga.edu/emt668/EMAT6680.2002/Curro/EMAT%206700/Chapter%206/Section2.html